OpenStudy (explicitvogue):

What is a polynomial function in standard form with zeros 1, -1, 2, and 4? A. y=x^4-6x^3+7x^2+6x-8 B. y=x^4-6x^3+9x^2+6x-8 C. y=x^3+7x^2+6x-8 D. x^4-x^3+x^2+x-8 I actually want help with how to solve this, I don't just want the answer. Fan + Medal to whoever can help the best :)

OpenStudy (explicitvogue):

@Michele_Laino could you help with this one? xx

OpenStudy (explicitvogue):

@pooja195

OpenStudy (michele_laino):

hint: the requested polynomial function have to contain such factors: \[\Large x - 1,\quad x + 1,\quad x - 2,\quad x + 4\]

OpenStudy (michele_laino):

so, please do this multiplication: \[\Large \left( {x - 1} \right) \cdot \left( {x + 1} \right) \cdot \left( {x - 2} \right) \cdot \left( {x + 4} \right) = ...?\] what do you get?

OpenStudy (explicitvogue):

I'm still confused, i dont understand alegbra at all

OpenStudy (explicitvogue):

i figured it out, it was A

OpenStudy (michele_laino):

I meant that the requested polynomial function can be written as below: \[\Large y = \left( {x - 1} \right) \cdot \left( {x + 1} \right) \cdot \left( {x - 2} \right) \cdot \left( {x - 4} \right)\]

OpenStudy (michele_laino):

Now, if we compute the indicated products, we get: \[\Large \begin{gathered} y = \left( {x - 1} \right) \cdot \left( {x + 1} \right) \cdot \left( {x - 2} \right) \cdot \left( {x - 4} \right) = \hfill \\ \hfill \\ = \left( {{x^2} - 1} \right) \cdot \left( {{x^2} - 6x + 8} \right) = \hfill \\ \hfill \\ = {x^4} - 6{x^3} + 7{x^2} - 6x - 8 \hfill \\ \end{gathered} \] so, your answer is right! @explicitvogue